# Dual-Branch Convolutional Framework for Spatial and Frequency-Based Image Forgery Detection

### v1

@@ 2 Literature Review
The detection of digital image tampering has been a long-standing area of research in digital forensics [1], gaining prominence as editing tools have become more sophisticated and accessible. The early methods in this field relied largely on handcrafted statistical and signal-processing techniques [2]. These methods aimed to expose inconsistencies introduced during the editing process, such as abrupt changes in lighting, irregular noise patterns, or compression artifacts. Techniques like error level analysis (ELA) [3], discrete cosine transform (DCT) [4] coefficient inspection, and sensor pattern noise analysis [5] were developed to identify these inconsistently manipulated areas. They laid the foundation for automated detection pipelines and demonstrated the feasibility of identifying image tampering using statistical measures. However, their effectiveness was often limited to specific types of edits, making them struggle to generalize across diverse image sources and varying post-processing conditions. [6]
@@
As the field progressed, tampering detection methods diverged into two broad categories: passive (or blind) and active approaches. Passive methods operate exclusively on the received images, relying on statistical, structural, or perceptual inconsistencies to locate manipulated areas [7]. These approaches gained widespread adoption due to their ability to work with any digital image, regardless of its source, making them ideal for forensics applications across platforms and media. In contrast, active methods embed digital watermarks, hash codes, or unique identifiers at the time of image capture [8]. These embedded marks enable verification and tamper detection with a high degree of reliability. However, active methods require access to or control of the original camera or editing process, making them impractical for situations where images are circulated online or captured by arbitrary devices. As a result, passive methods have emerged as the dominant approach for digital tampering detection.
@@
With advances in machine learning and deep learning, tampering detection has entered a new era. Convolutional neural networks (CNNs) [9, 10] and, more recently, transformers [11] have demonstrated remarkable capabilities for learning hierarchical, context-aware features that expose inconsistencies introduced during editing. These deep learning methods have proven highly effective for challenging tampering scenarios, including splicing, copy-move, inpainting, and deepfake generation. Models such as multi-task fully convolutional networks have demonstrated strong performance in locating tampered areas by focusing on noise inconsistencies and spatial artifacts. Meanwhile, noiseprint-based approaches [12] have leveraged deep feature embeddings to isolate camera or sensor characteristics, making it possible to pinpoint even highly sophisticated manipulations. Newer methods utilize attention mechanisms, self-supervised learning, and multi-scale feature aggregation to further improve detection and generalization across diverse datasets and editing techniques [13, 14].
@@
Despite significant advances, the field still faces a range of critical challenges. The emergence of highly realistic, AI-generated images has pushed the boundaries of traditional detection pipelines [15], making it increasingly difficult to distinguish between authentic and manipulated media. The scarcity of large-scale, high-quality, and well-annotated tampering datasets also limits the ability of deep learning models to learn robust and generalizable representations. In addition, the vulnerability of deep detection methods to adversarial attacks presents a serious concern, as carefully crafted perturbations can fool state-of-the-art detection systems [16]. Another challenge lies in extending detection methods to multi-modal settings, such as verifying the consistency between images and associated text, or detecting inconsistencies across frames in a video [17, 18, 19].

### v2

@@ 2 Literature Review
The detection of digital image tampering has been a long-standing area of research in digital forensics [mishra2013digitalimagetamperdetection], gaining prominence as editing tools have become more sophisticated and accessible. The early methods in this field relied largely on handcrafted statistical and signal-processing techniques [Christlein_2012]. These methods aimed to expose inconsistencies introduced during the editing process, such as abrupt changes in lighting, irregular noise patterns, or compression artifacts. Techniques like error level analysis (ELA) [krawetz2007picture], discrete cosine transform (DCT) [wandji2013detectioncopymoveforgerydigital] coefficient inspection, and sensor pattern noise analysis [matthews2019rethinkingimagesensornoise] were developed to identify these inconsistently manipulated areas. They laid the foundation for automated detection pipelines and demonstrated the feasibility of identifying image tampering using statistical measures. However, their effectiveness was often limited to specific types of edits, making them struggle to generalize across diverse image sources and varying post-processing conditions. [thakur2020recent]
@@
As the field progressed, tampering detection methods diverged into two broad categories: passive (or blind) and active approaches. Passive methods operate exclusively on the received images, relying on statistical, structural, or perceptual inconsistencies to locate manipulated areas [LIN201829]. These approaches gained widespread adoption due to their ability to work with any digital image, regardless of its source, making them ideal for forensics applications across platforms and media. In contrast, active methods embed digital watermarks, hash codes, or unique identifiers at the time of image capture [article]. These embedded marks enable verification and tamper detection with a high degree of reliability. However, active methods require access to or control of the original camera or editing process, making them impractical for situations where images are circulated online or captured by arbitrary devices. As a result, passive methods have emerged as the dominant approach for digital tampering detection.
@@
With advances in machine learning and deep learning, tampering detection has entered a new era. Convolutional neural networks (CNNs) [liu2022dunetdualencoderunetimage, katiyar2022imageforgerydetectioninterpretability] and, more recently, transformers [wang2024timelysurveyvisiontransformer] have demonstrated remarkable capabilities for learning hierarchical, context-aware features that expose inconsistencies introduced during editing. These deep learning methods have proven highly effective for challenging tampering scenarios, including splicing, copy-move, inpainting, and deepfake generation. Models such as multi-task fully convolutional networks have demonstrated strong performance in locating tampered areas by focusing on noise inconsistencies and spatial artifacts. Meanwhile, noiseprint-based approaches [cozzolino2018noiseprintcnnbasedcameramodel] have leveraged deep feature embeddings to isolate camera or sensor characteristics, making it possible to pinpoint even highly sophisticated manipulations. Newer methods utilize attention mechanisms, self-supervised learning, and multi-scale feature aggregation to further improve detection and generalization across diverse datasets and editing techniques [hao2021transforensicsimageforgerylocalization, liu2022multiscalewavelettransformerface].
@@
Despite significant advances, the field still faces a range of critical challenges. The emergence of highly realistic, AI-generated images has pushed the boundaries of traditional detection pipelines [pei2024deepfakegenerationdetectionbenchmark], making it increasingly difficult to distinguish between authentic and manipulated media. The scarcity of large-scale, high-quality, and well-annotated tampering datasets also limits the ability of deep learning models to learn robust and generalizable representations. In addition, the vulnerability of deep detection methods to adversarial attacks presents a serious concern, as carefully crafted perturbations can fool state-of-the-art detection systems [gragnaniello2018analysisadversarialattackscnnbased]. Another challenge lies in extending detection methods to multi-modal settings, such as verifying the consistency between images and associated text, or detecting inconsistencies across frames in a video [song2024trinitydetectortextassistedattentionmechanisms, song2025learningmultimodalforgeryrepresentation, liu2025forgerygptmultimodallargelanguage].